# Group Testing with General Correlation Using Hypergraphs

### v1

@@ 9 Conclusion and Future Work
In this paper, we developed a novel framework for group testing with general statistical correlation among nodes. Our model considers a probability distribution on all possible candidate subsets of nodes, captured by a hypergraph. We developed group testing algorithms that exploit correlation in two settings, namely adaptive and semi-non-adaptive group testing. In the adaptive setting, our algorithm consists of two stages: informative testing and individual testing. We demonstrate that the number of tests required is a function of the entropy and the expected number of infections, achieving order-wise optimality when the expected number of infections is less than the entropy. In particular, when the number of infections is concentrated around its mean, the number of required tests is upper bounded by $O(H(X)+\mu)$, in expectation. We provided several examples illustrating both the worst-case optimality and limitations of the algorithm. The algorithm remains robust in noisy settings (with symmetric noise), as it can correctly update its belief despite errors.
We further introduced and studied a semi-non-adaptive setting– a model in between adaptive and non-adaptive settings. In this model. tests are designed independently from prior test results (similar to non-adaptive group testing), but it is assumed that testing can stop once the infected set is found (sequentially as a function of prior test results). In this setting, we provided a group testing strategy that requires at most $O(\mu H(X)+\mu\log n)$, in expectation. The algorithm can be extended to the noisy group testing setting with symmetric noise.
@@ Appendix A Modeling Statistical Models with Hypergraphs
Appendix B $\Omega(n)$ Tests are Needed in Example 6.1
Theorem B.1.
Any algorithm with a constant probability of success with the graph in Example 6.1 requires at least $\Omega(n)$ tests.
Proof.
First, note that every test of size two or greater will yield positive results, rendering them redundant. Consequently, the problem is effectively simplified to individual testing when a single individual is found to be negative.
Every test is positive unless we find the negative node, so it takes $\Omega(n)$ tests to identify that particular node if we want to have a constant probability of success.
Appendix C A Proof for Example 6.2
Consider all the subsets of $S$ where $|S|=d+1$. We prove that the average number $E(S)$ is $c(d+1)$. There are (${n\choose d+1}$ of subsets of size $d+1$, and each edge appears in $n-d$ of such subsets. Hence the average number of edges inside $S$ is $\frac{(n-d)|E|}{{n\choose d}}=\frac{c(n-d){n\choose d}}{{n\choose d+1}}=c(d+1)$.
Now if the algorithm has additional knowledge that the target set is in a subset $S$ where $|S|=d+1$ and $E(S)\geq c(d+1)$, it means there are $c(d+1)$ candidates for the single noninfected node, hence any algorithm needs to perform $O(d)$ tests for a constant probability of success.
Appendix D Proof of [ACÖ23] From Section 5
@@
Appendix E Proof of Theorem 7.1
We already proved the theorem when every edge has size $d$. We are using the same algorithm as described for the $d-$regular case, where each node is in the test with probability $1/\mu$ at random, and slightly change the analysis. We first prove that when the target edge is in $i$’th subgraph, $\mu logm_{i}$ tests are enough to detect it. The argument is very similar to [GLS22]. Let $\ell=\mu-c$ a and $u=\mu+c$ where $c=o(\mu)$ and $\Pr_{e\sim\mathcal{D}}(\ell<|e|<u)=1-o(1)$. We only focus on the case when $\ell<|e|<u$. For any pair of edges $e_{1}$ and $e_{2}$ in $G_{i}$, we call a test $T$ separates them if $T$ contains at least a node in $e_{1}$ and no node in $e_{2}$, or $T$ contains at least a node in $e_{2}$ and no node in $e_{1}$. The probability that a random test separates $e_{1}$ and $e_{2}$ is at least $\frac{(1-1/\mu)^{u}}{\mu}=\Theta(\frac{1}{\mu})$. There are (${m_{i}\choose 2}=O(m_{i}^{2})$ different pairs, so if we do $3\mu\log m_{i}$ tests, by union bound, we can separate any two edges with high probability. This in turn means that we can rule out all the edges that are not the target edge. Hence, by knowing that the target edge is in $i$’th subgraph we can recover it using $\mu\log m_{i}$ tests. Using a similar argument, in expectation we perform $\mu H(X)$ tests in expectation when we know the subgraph that contains the target edge.
Now we perform random testing until there is at least one subgraph with exactly one edge and we perform $2\mu\log n$ additional test to verify it. Using a similar argument as above, the probability that the target edge remains after the additional tests is $O(1/n^{2})$ and there are at most $n$ subgraphs, hence the probability that we output it at any point is at most $O(1/n)$. This completes the proof

### v2

@@ 9 Conclusion and Future Work
In this paper, we developed a novel framework for group testing with general statistical correlation among nodes. Our model considers a probability distribution on all possible candidate subsets of nodes, captured by a hypergraph. We developed group testing algorithms that exploit correlation in a variety of settings, namely adaptive, semi-non-adaptive group testing, noisy etc. We provide constructive testing algorithms and upper bound the expected number of tests they need to attain asymptotically vanishing probabilities of error. We investigate special cases of interest, such as random regular hypergraphs, and fine tune our algorithms therein to obtain order-wise optimal number of tests for vanishing probabilities of error for sparse and dense graphs. We also obtain lower bounds on the expected number of tests needed for attaining vanishing probability of error in the general case and several special cases, which in turn reveal if the expected number of tests for our constructive algorithms can be reduced through improved design. The modifications of the algorithms we provide have several desirable characteristics: 1) ability to schedule tests in advance without waiting for results 2) innate robustness to errors in estimation of parameters 3) tradeoffs between number of tests and accuracy of estimates and accuracy of tests.
@@
10 Acknowledgment
The authors would like to acknwledge Dominic Olaguera-Delogu for helpful discussions and contributions to Example 6.7. The work of Hesam Nikpey and Shirin Saeedi Bidokhti was supported by the NSF CAREER award 2047482. The work of Saswati Sarkar was supported by NSF grant 2008284, and SIGA. The authors gratefully acknowledge this support.
@@ Appendix A Modeling Statistical Models with Hypergraphs
Appendix B Proof of [ACÖ23] From Section 5
@@
Appendix C Proof of Theorem 7.1
We already proved the theorem when every edge has size $d$. The assumption in the theorem implies that if $e^{*}$ is the target edge, $\Pr(|e^{*}|<u)=1-o(1/n)$. We are using the same random testing strategy as for the $d-$regular case, with the difference that now each node is selected in any given test with probability $1/u$ (instead of $1/d$). The algorithm is the same as well, except that a preprocessing step is added in the beginning in which any edge $e$ whose size is at least $u$ is removed. Note that no node is removed in the preprocessing.
The only ways the algorithm does not return a target edge is when at least one of the following events happen: 1) the target subgraph does not enter $SG$ by $2u(n-5\log n)$ tests 2) at least one non-target subgraph that enters $SG$ before or at the same test as the target subgraph remains candidate for $10u\log n$ tests, 3) the number of infected nodes is at least $u$ (since all the infected edges together constitute the target edge, then the target edge is removed in the preprocessing). Note that the first event implies that either the target subgraph does not become a candidate by the end of all tests or $10u\log n$ tests are not conducted between when it becomes candidate and all tests end. We refer to these three events as $err_{1},err_{2},err_{3}$ and show that the respective probabilities of $err_{1}\cap(err_{3})^{c},err_{2}\cap(err_{3})^{c},err_{3}$, $\delta_{n}^{1},\delta_{n}^{2},\delta_{n}^{3}$ are $o(1/n).$ Then the probability of the union of the three error events is also $o(1/n)$ by union bound.
For upper-bounding $\delta_{n}^{1},\delta_{n}^{2}$ we first upper bound the probability that an edge $e$ that is not the target edge survives $T$ successive random tests given that $I<u.$ Suppose $e^{*}$ is the target edge. Note that if $e\not\subset e^{*}$, then $e$ must have at least one negative node and it is detected as negative if it is selected in a test and none of the positive nodes are selected for that test. If $e\subset e^{*}$, then there is a node in $e^{*}$ that is not in $e$ and this node is positive. Call the node in each of the cases as $v$. Note that if $v$ appears in a test without any other node in $e^{*}$, then $e$ will be eliminated in both cases (In the second case, $v$ just has to appear in a test without any node in $e$, but $e\subseteq e^{*}$; so if $v$ appears without any other node in $e^{*}$, it also appears without any other node in $e.$). Given that there are at most $u$ positive nodes overall, the probability that $v$ appears in a test without other nodes in $e^{*}$ is at least $(1/u)(1-(1/u))^{u}.$ Thus, given that $I<u$, probability that $e$ is ruled out as a target edge in a random test is at least $(1/u)(1-(1/u))^{u}.$ The conditional probability that $e$ survives after $T$ tests is therefore at most $(1-(1/u)(1-(1/u))^{u})^{T}\leq e^{-(1-1/u)^{u}T/u}$ since $1-x\leq e^{-x}$ with $(1/u)(1-(1/u))^{u}$ as $x.$ Now, noting that $(1-1/u)^{u}\geq 1/4$, for all $u\geq 2$, we have the conditional probability upper bounded by $e^{-T/(4u)}.$ Since the probability that $I<u$ is at most $1$, the probability that edge $e$ that is not the target edge survives $T$ successive random tests and $I<u$ is at most $e^{-T/(4u)}.$
Now, we upper bound the probability of the intersection of $err_{1}$ and that $I<u$ given that $G_{i}$ is the target subgraph. First note that $m_{i}\leq 2^{n}$ for each $i.$ Thus, $u\log m_{i}\leq un$, and $u\log m_{i}+8u\log n<2u(n-5\log n)$ for all large $n$. Thus, 1) the algorithm conducts at least $u\log m_{i}+8u\log n$ tests, and 2) a necessary condition for $err_{1}$ to happen is that one or more edges other than the target edge in $G_{i}$, survive $u\log m_{i}+8u\log n$ tests. The probability that a given edge $e$ in $G_{i}$ that is not the target edge survives $u\log m_{i}+8u\log n$ successive random tests and $I<u$ is at most $e^{-(\log m_{i}+8\log n)/4}=(1/m_{i})(1/(n^{2})).$ Now, there are $m_{i}-1$ edges in $G_{i}$ that are not target edges. Thus, by union bound, the conditional probability is at most $1/(n^{2}).$ Since the upper bound does not depend on $i$, this is also the upper bound on $\delta_{n}^{1}$, the probability of the intersection of $err_{1}$ and $I<u.$
We now upper bound the probability of the intersection of $err_{2}$ and $I<u.$ The probability that edge $e$ that is not the target edge survives $10u\log n$ successive random tests and $I<u$ is at most $e^{-10(u\log n)/(4u)}=e^{-2.5\log n}=(1/n^{2.5}).$ At most $n\log n$ non-target subgraphs enter $SG$ on or before the target subgraph. So the probability that $I<u$ and at least one of those remain candidate after $10u\log n$ tests is at most the probability that $I<u$ and at least one among $n\log n$ target edges survive $10u\log n$ tests. By union bound the probability that $I<u$ and at least one among $n\log n$ target edges survive $10u\log n$ tests is at most $(n\log n)/n^{2.5}=(\log n)/n^{1.5}.$ This also upper bounds $\delta_{n}^{2}$, the probability of the intersection of $err_{2}$ and $I<u.$
By assumption in the theorem, $\delta_{n}^{3}$ is $o(1/n).$
We now analyze the expected number of tests, similar to that for $d$-regular graphs. Note that with probability $1-o(1/n)$ neither of $err_{1},err_{2}$, or $err_{3}$ happen, and then when additionally the $i$’th subgraph is the target subgraph, the algorithm performs $u\log m_{i}+8u\log n+10u\log n$ tests. Note that when any of the $err_{i}$’s happens, at most $2un$ tests are performed. Hence the expected number of tests is at most:
$\sum_{i}\mathbb{P}(G_{i})[(1-(\delta_{n}^{1}+\delta_{n}^{2}+\delta_{n}^{3}))(u\log(m_{i})+8u\log n+10u\log n)+(\delta_{n}^{1}+\delta_{n}^{2}+\delta_{n}^{3})2un]$
$\leq u\sum_{i}\left[(\sum_{e\in G_{i}}p(e))\log(m_{i})\right]+18u\log n+o(u).$
Recall that $m_{i}\leq 2/p(e)$ for all $e\in E_{i}$. Hence the RHS is
$\leq u\sum_{i}\sum_{e\in G_{i}}p(e)\log(2/p(e))+18u\log n+o(u)$
$=u\sum_{e\in E}p(e)(1+\log(1/p(e)))+18u\log n+o(u)$
$=u+uH(X)+18u\log n+o(u).$ (11)
And hence the total number of tests is at most $O(uH(X)+u\log n)$ in expectation.